An empty graduated cylinder weighs 75g. When filled with 50ml of an unknown liquid it weighs 125g.  What is the density of the liquid?

a. 1.00 g/mL b. 1.67 g/mL c. 2.50 g/mL d. 4.00 g/mL




1
Expert's answer
2021-04-15T10:32:25-0400

Since the mass of the empty cylinder is 75g, and the mass of the filled cylinder is 125g, the mass of the liquid is:


"m = 125g-75g = 50g"

Let "V = 50ml" be the volume of the liquid. By definition, the density of any substance is:


"\rho = \dfrac{m}{V}"

Substituting the values, obtain:


"\rho = \dfrac{50g}{50ml} = 1\space g/ml"

Answer. a. 1.00 g/mL.